OpenStudy (anonymous):

\[\tan(\frac{ 1 }{ 2 }\cos ^{-1}\frac{ \sqrt{5} }{ 3 }) = ?\]

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73 @satellite73

OpenStudy (anonymous):

half angle formula for this one

OpenStudy (anonymous):

\[\tan(\frac{\theta}{2})=\frac{\sin(\theta)}{1+\cos(\theta)}\] if i remember correctly

OpenStudy (anonymous):

since \(\cos^{-1}(\theta)=\frac{\sqrt{5}}{3}\) this is \[\frac{\sin(\theta)}{1+\frac{\sqrt{5}}{3}}\]

OpenStudy (anonymous):

you do not need \(\theta\) you only need \(\cos(\theta)\) and \(\sin(\theta)\)

OpenStudy (anonymous):

to find \(\sin(\theta)\) is the same as being asked "what is \(\sin(\theta)\) if \(\cos(\theta)=\frac{\sqrt{5}}{3}\) ? we only need the other side of the triangle, which we get by pythagoras

OpenStudy (anonymous):



OpenStudy (anonymous):

\[?=\sqrt{3^2-\sqrt{5}^2}=\sqrt{9-5}=\sqrt{4}=2\] so \[\sin(\theta)=\frac{2}{3}\]

OpenStudy (anonymous):

penultimate answer is \[\frac{\frac{2}{3}}{1+\frac{\sqrt{5}}{3}}\] then you can clean this up

OpenStudy (anonymous):

@satellite73 my Approach: cos^-1 sqrt5/3 = t cos t = sqrt5/3 tan( t/2) = (Sint/2)/cost/2 1-2sin^2 t/2 = cost 2cos^2 t/2 = cos t

OpenStudy (anonymous):

Cos t/2 = sqrt(sqrt5 + 3)/sqrt6 Sin t/2 = sqrt(3-sqrt5)/sqrt6

OpenStudy (anonymous):

@satellite73 R u there:)

OpenStudy (anonymous):

i am here

OpenStudy (anonymous):

Can u Check My Solution..

OpenStudy (anonymous):

and @satellite73 i am Unable to PM u....Plzz Make it Liable

OpenStudy (anonymous):

i don't think your half angle formula for cosine is right, let me check

OpenStudy (anonymous):

\[\cos(t)=2\cos^2(\frac{t}{2})-1\] i think

OpenStudy (anonymous):

much easier to use the half angle formula for tangent, which contains sine and cosine

OpenStudy (anonymous):

Is nt it tan t/2 = (1-cos t) / sin t

OpenStudy (anonymous):

either way it is the same

OpenStudy (anonymous):

you can use \[\frac{1-\cos(t)}{\sin(t)}=\frac{\sin(t)}{1+\cos(t)}\]

OpenStudy (anonymous):

it is easy to see that they are equal, cross multiply and you get \[\sin^2(t)=1-\cos^2(t)\] which we know is true

OpenStudy (anonymous):

Yup....thxxxx....